[Ncep.nhc.nco_contacts] TSB Morning Rounds - Tue Dec 06, 2016

TSB Admin nhc.tsbadmin at noaa.gov
Tue Dec 6 14:17:54 UTC 2016


Here is today's summary of NHC computer operations:
--- Craig Mattocks

----------

TAFB:
1. No TAFB text products had been updated on the NHC website since  
yesterday morning. TSB Developer Dave Zelinksy noticed this morning that  
the LDM running on the logger socket ingest system had become partially  
dysfunctional. While the socket software was working fine and the LDM was  
running (products were being ingested), the LDM wsn't sending those  
products downstream and ldmadmin wasn't responding to any stop or restart  
commands (it was just spewing endless "Waiting for the LDM to terminate..."  
messages). A Priority-1 NCO Help Desk ticket was submitted and NCO Linux  
SysAdmin Curt Steinmetz rebooted logger. This restored the LDM to full  
functionality.
HSU:
No new issues.
Other items:

NHC RFCs:


-------------- next part --------------
An HTML attachment was scrubbed...
URL: https://www.lstsrv.ncep.noaa.gov/pipermail/ncep.nhc.nco_contacts/attachments/20161206/78b7dffe/attachment.html 


More information about the Ncep.nhc.nco_contacts mailing list